OpenStudy (anonymous):

How do I prove this: If Sn = (1+1/n)^n, then (sn) is strictly increasing. I know the defn to be strictly increasing is: A sequence is strictly increasing if and only if for all n, Sn < Sn+1

OpenStudy (anonymous):

A proof by induction is one way to go about it. First show it holds for some base case, say \(n=1\): \[s_n=\left(1+\frac{1}{n}\right)^n\\ s_1=2\\ s_2=\frac{9}{4}=2.25\] \(s_1<s_2\), so the base case is true. Now assume the claim holds for \(n=k\), i.e. that \(s_k<s_{k+1}\). You must now show that this implies \(s_{k+1}<s_{k+2}\). Under the assumption, you have \[\left(1+\frac{1}{k}\right)^k<\left(1+\frac{1}{k+1}\right)^{k+1}\] You know that \[s_{k+2}=\left(1+\frac{1}{k+2}\right)^{k+2}=\left(1+\frac{1}{k+2}\right)\left(1+\frac{1}{k+2}\right)^{k+1}\] Now what you want to do is compare this with \(s_{k+1}\). Here is what you need to establish: \[\underbrace{\color{red}{\left(1+\frac{1}{k+1}\right)^{k+1}}}_{\large s_{k+1}}<\underbrace{\left(1+\frac{1}{k+2}\right)\color{red}{\left(1+\frac{1}{k+2}\right)^{k+1}}}_{\large s_{k+2}}\] First compare the red quantities. Since they're raised to the same power, you can focus on what's inside the parentheses. As \(k\) gets bigger, \(1+\dfrac{1}{k+1}\approx1+\dfrac{1}{k+2}\). On the right side, the \(1+\dfrac{1}{k+2}\) factor is greater than 1, which will constantly increase the value of the right side. This confirms that the right side will always be greater, thus completing the proof.